# LEARNING TO WALK ON NEW GROUND:

### v1

@@ INTRODUCTION
A brain-computer interface (BCI) records the user’s brain activity and converts these into computer commands, offering an alternative output channel that does not rely on muscular activity. Electroencephalography (EEG) is commonly used to record brain activity due to its affordability, practicality, and non-invasiveness. The primary application of BCIs lies in restoring lost control, particularly in communication. One notable example is the visual BCI speller, where users can select symbols by focusing their gaze on them when displayed on a screen. This technology proves invaluable for individuals facing challenges such as amyotrophic lateral sclerosis (ALS), where progressive loss of voluntary motor control makes speaking and typing difficult [1].
@@
A major challenge for a BCI is that prior to its usage, a machine learning model capable of classifying unseen brain signals needs to be calibrated on labeled EEG data from the same user. This is necessary because users display individual patterns of brain activity (subject-to-subject variability). Additionally, the same user might show different patterns over multiple days of use (session-to-session variability) and even within-session non-stationarity. To mitigate any negative effects of these confounders, the user is guided through an initial stage to record brain activity while being instructed which symbol to attend to.
@@
While a trained classification model is necessary to practically use a BCI for the intended application, the calibration recording delays the immediate deployment of the BCI and may be prohibitive specifically for users with a limited attention span. In general, the necessity of calibration may impede the acceptance and widespread adoption of the BCI system by patients as well as healthy users. Encouragingly, recent advancements in BCI technology have surfaced which offer potential solutions to alleviate this challenge.
@@
The first advancement involves selecting an informative brain signal feature to minimize the duration of the calibration phase. BCIs can be driven by various brain signals, often evoked by advanced stimulus protocols. Many popular stimulus protocols induce one or more event-related potentials (ERPs). Among these, the visual evoked potential (VEP), triggered by a flash, stands out. VEP-based BCIs are widely embraced due to their effectiveness across diverse user populations [2]. The VEP can be effectively used in three different ways [3].
@@
Firstly, in a time-modulated (t-VEP) based BCI, stimuli are sequentially presented to reduce temporal overlap, resulting in a relatively slow paradigm. Secondly, in a frequency-modulated (f-VEP) based BCI, each stimulus simultaneously and rapidly flashes at a unique frequency and phase [4]. Despite its speed, f-VEP faces limitations due to the restricted range of narrow-band options and potential artefacts that may obscure signals. Thirdly, in a code-modulated (c-VEP) based BCI, each stimulus rapidly flashes with a pseudo-random noise-code [5]. In this protocol, stimulus sequences are optimized to be dissimilar, ensuring that their corresponding brain activity is dissimilar as well. Remarkably, c-VEP BCI has recently demonstrated unprecedented performance [6].
@@
The second advancement involves the choice of the decoding approach. In the case of c-VEP, a method was developed, termed ‘reconvolution’, which relies on a forward model embedded in a canonical correlation analysis (CCA) [7]. This model characterizes the response to a sequence of flashes as the linear summation of responses to individual flashes. This reconvolution approach substantially reduced the number of trainable parameters while it simultaneously increased the number of samples that were available as training data points. This not only decreased the required training data but also empowered the model to predict responses to unseen stimulus sequences. Recently, this reconvolution CCA method was shown to achieve remarkable performances on c-VEP data event without the need for a calibration session, by finding the stimulus sequence that best fits the data in a trial [8].
@@
In a more recent development, a novel classification approach was introduced for ERP based BCI. The approach, termed ‘unsupervised mean-difference maximization’ (UMM) does not require labeled calibration data [9]. UMM does not act on every single epoch, i.e., the evoked response of a single stimulus, but on a set of epochs belonging to one control command, e.g., one trial that leads to the selection of a symbol in a spelling application. Therefore, UMM has similarities to the aforementioned CCA method. While the objective of CCA is to maximize the explained variance, UMM identifies an attended target symbol by maximizing the distance between target and non-target ERPs belonging to one trial. Diverging from the CCA method, UMM incorporates several regularization approaches, including the use of block-Toeplitz covariance matrices [10] for determining domain-specific distances, and it can take advantage of a built-in confidence metric. Collectively, UMM has demonstrated impressive performance across various ERP datasets, without the need for a training session [9].
@@
Dataset: We assessed the efficacy of the CCA and UMM zero-training approaches using an open-access c-VEP dataset [11]. Comprehensive details about this dataset can be found in the original study [8]. For the current study, we only used the part of this dataset labeled as ‘offline experiment’, in which 30 participants engaged in a copy-spelling task. EEG data were recorded from 8 electrodes placed following the 10-10 system (Fz, T7, O1, POz, Oz, Iz, O2, T8) amplified using a Biosemi Active2 amplifier and sampled at a frequency of 512 Hz.
Throughout the experiment, participants interacted with a $4\times 5$ matrix speller displayed on a 12.9 in iPad Pro with a 60 Hz refresh rate and a $1920\times 1080$ px resolution. The $N=20$ cells within this matrix each measured 3.1 cm $\times$ 2.8 cm, with a 0.4 cm separation both horizontally and vertically between cells. The cells were presented against a mean-luminance gray background. Each cell $i\in\{1,\dots,N\}$ was luminance modulated using a unique binary stimulus sequence at full contrast, with a 1 encoding a white cell and a 0 a black cell. The stimulus sequences were carefully chosen from an optimized subset of Gold codes [12]. These sequences were modulated such that they contained flashes of only two durations: a short flash of 16.67 ms and a long flash of 33.33 ms. The sequences had a length of $126=2*(2^{6}-1)$ bits and at 60 Hz cycling through a stimulation code once took 2.1 s.
@@
Alternatively to this instantaneous CCA, CCA can learn across trials. Specifically, previous trials can be included to improve the estimates for the current trial, as described in [8]. Equation 1 can be formulated using the spatio-temporal cross-covariance ${\bm{\Sigma}}_{\mathbf{X}\mathbf{M}_{i}}\in\mathbb{R}^{C\times M}$, the spatial covariance ${\bm{\Sigma}}_{\mathbf{X}}\in\mathbb{R}^{C\times C}$ and the temporal covariance ${\bm{\Sigma}}_{\mathbf{M}_{i}}$:
@@
In summary, in this work, we applied two versions of CCA. For both we used the empirical covariance matrix identical to the original work [8]. The first version, denoted CCA_e1, was instantaneous and estimated the covariance from the current trial only. The second version, denoted CCA_ec, was cumulative and used previous trials for covariance estimation to facilitate decoding of the current trial.
Code for the CCA approach is available at: https://github.com/thijor/pyntbci.
@@
Instead of determining the attended symbol of the current trial by maximizing the Euclidean distance between flash and non-flash ERPs across all $N$-many hypotheses, the metric is first normalized using the inverse of the global feature covariance matrix ${\bm{\Sigma}}\in\mathbb{R}^{D\times D}$ in order to better cope with non-spherical feature distributions in the feature space $\mathbb{R}^{D}$, known as the Mahalanobis distance:
@@
Please note that the covariance matrix can be estimated based on the epochs of the current trial only, which makes UMM an instantaneous decoding approach that does not require calibration data. Due to the challenging ratio of the feature dimensionality and the number of epochs contained in a single trial, we used a block-Toeplitz regularization with tapering to obtain a more robust estimate of the covariance matrix [10].
Alternatively to this instantaneous use of UMM, knowledge from previous trials about both, the estimated class means and the covariance matrix can be included for obtaining improved estimates for the current trial, as described in [9]. Specifically, as the covariance matrix can be calculated without label information the covariance matrix can more robustly be estimated by also using epochs from previous trials [9]. A similar approach can be used to more robustly estimate the flash and non-flash ERPs, by using information from previous trials. However, for this, label information is required. UMM simply uses its own predictions from previous trials as pseudo labels (sometimes also referred to as naive labeling). This approach is made more robust, by weighting the mean estimates from previous trials by UMM’s confidence in these previous trials [9]. Specifically, if UMM is very certain of its own prediction, these ERP means will have more weight in later trials and vice versa.
In summary, for this work, we used two versions of UMM. For both versions we used the block-Toeplitz regularized covariance matrix. The first version, denoted UMM_t11, is instantaneous by estimating the covariance and means only from the current trial. The second version, denoted UMM_tcw, is cumulative by using previous trials to facilitate decoding of the current trial.
Code for the UMM approach is available at: https://github.com/jsosulski/umm_demo.
@@ RESULTS
The instantaneous CCA method (CCA_e1) achieved a performance of 0.06 at 1.05 s, while the instanteneous UMM (UMM_t11) reached 0.09. At this time point, UMM_t11 significantly outperformed CCA_e1. Instead, at any further time point CCA_e1 significantly surpassed the accuracy of UMM_t11. At the maximum trial length of 31.5 s, CCA_e1 reached a classification of 0.96 which was significantly higher than UMM_t11 with 0.89.
The cumulative versions always outperformed the instantaneous version for both CCA and UMM. Notably, the instantaneous CCA (CCA_e1) performed almost on par with its cumulative version (CCA_ec) for trial durations longer than 25.2 s. Specifically, at 27.3 s, CCA_e1 reached an accuracy of 0.95 while CCA_ec reached 0.97, which was not significantly higher. Also at the 29.4 and 31.5 s trial durations, CCA_ec did not significantly surpass CCA_e1.
Figure 1: Bandpass hyper-parameters for CCA and UMM. Depicted are the grand average classification accuracy for CCA and UMM across varying highpass (left) and lowpass (right) cutoff values. Here, a single-trial duration of 31.5 s is used. When varying the highpass, the lowpass remained at 40 Hz, and when varying the lowpass, the highpass remained at 6 Hz. In order, the symbols behind a method refer to: the type of covariance matrix being empirical (e) or block-Toeplitz (t); covariance matrices computed instantaneously (1) or cumulative (c); and the mean vectors (of UMM) computed instantaneously (1) or using a weighted cumulative average (w). The dashed horizontal gray line depicts theoretical chance level decoding (5%).
Figure 2: Decoding curve for CCA and UMM. Depicted are the grand average classification accuracy for CCA and UMM across varying single-trial durations. Here, a bandpass of 6-50 Hz is used. In order, the symbols behind a method refer to: the type of covariance matrix being empirical (e) or block-Toeplitz (t); covariance matrices computed instantaneously (1) or cumulative (c); and the mean vectors (of UMM) computed instantaneously (1) or using a weighted cumulative average (w). The vertical grid lines indicate when a new code cycle starts (every 2.1 s). The dashed gray horizontal line depicts theoretical chance level decoding (5%).
@@ DISCUSSION
This study introduced UMM to c-VEP BCI. UMM is a calibration-free method that originates from the ERP domain. We conducted a systematic comparison with the conventional c-VEP method employing CCA. Both methods underwent evaluation in an instantaneous manner, classifying each single-trial without prior knowledge, as well as in a cumulative way, utilizing previously classified trials as training data. The ultimate goal was to establish a c-VEP BCI with enhanced usability and broader adoption potential by eliminating the requirement for a tedious initial calibration session.
Both the conventional method, CCA, and the novel approach, UMM, hinge on the principle that selecting a single symbol from a set of candidate symbols is considerably simpler than tackling the more complex task of reconstructing the entire stimulus sequence. In this study, for example, the number of symbols was $N=20$, resulting in just 20 candidate stimulus sequences for evaluation. In contrast, the exhaustive reconstruction of the stimulus involves an exponential growth, reaching already $2^{60}$ potential sequences for a 1-second stimulus at a presentation rate of 60 Hz.
@@
Furthermore, we demonstrated the capability of instantaneous classification, where CCA reached higher than 90 % accuracy at 14.70-second trials and UMM 89 % at 29.40 s. Instantaneous decoding is a very nice feature to have, as the classifier can remain fully flexible, automatically adapting to non-stationary data distributions.
Both CCA and UMM make different assumptions and exploit them. These assumptions may be met to different degrees by different datasets. For instance, CCA strongly leverages the sequential structure and large overlap between responses to adjacent stimuli in c-VEP datasets. Additionally, the current event definition used in this study focused on the target ERP alone, while discerning between ERPs associated with short and long flashes. Furthermore, UMM maintains its operations within the original EEG feature space, whereas CCA operates in the component space.
In contrast, UMM searches for the stimulus sequence with the largest target to non-target ERP distance, potentially rendering it less susceptible to slow drifts in the data, as evidenced by the highpass analysis. Besides, CCA uses the empirical covariance matrix, which can be challenging to accurately estimate with limited data, while UMM employs domain-specific regularization techniques such as shrinkage and a block-Toeplitz covariance matrix [10]. Lastly, in the UMM formulation, the mean estimates for the ERPs are improved using previously classified trials, not in an optimistic way assuming all previous trials were correctly classified, but instead carefully weighting these mean updates based on the confidence of each previous trial. By gaining a comprehensive understanding of the strengths and limitations of both methods, there exists the potential to develop refined versions tailored to specific characteristics of novel datasets through thoughtful hyper-parameterization.
It is essential to highlight that both CCA and UMM methods necessitate a specific stimulus protocol involving repetitions, and require knowledge about the precise timing and sequence of stimuli within a single trial (e.g., the selection of one symbol). While such information is typically available in BCI protocols using evoked responses, it may not seamlessly extend to other responses (e.g., induced responses like the sensorimotor rhythm). Moreover, these decoding methods are only applicable in benchmarks if such sequence information is provided, as demonstrated in MOABB [13]. This characteristic classifies the studied CCA and UMM methods as semi-supervised, given that stimulus information is requisite, while label information is not required.

### v2

# EXPLORING NEW TERRITORY: CALIBRATION-FREE DECODING FOR C-VEP BCI J. Thielen1, J. Sosulski2, M. Tangermann1

@@ INTRODUCTION
A brain-computer interface (BCI) records the user’s brain activity and converts these into computer commands, offering an alternative output channel that does not rely on muscular activity. Electroencephalography (EEG) is commonly used to record brain activity due to its affordability, practicality, and non-invasiveness. The primary application of BCIs lies in restoring lost control, particularly in communication. One notable example is the visual BCI speller, where users can select symbols by focusing their gaze on them when displayed on a screen. This technology proves invaluable for individuals facing challenges such as amyotrophic lateral sclerosis, where progressive loss of voluntary motor control makes speaking and typing difficult [verbaarschot2021].
@@
Prior to BCI usage, a machine learning model capable of classifying unseen brain signals needs to be calibrated on labelled EEG data from the same user, as individuals display different patterns of brain activity. Additionally, the same user might show different patterns over multiple days of use (session-to-session variability) and even within-session non-stationarity. To mitigate any negative effects of these confounders, the user is guided through an initial stage to record brain activity while being instructed which symbol to attend to.
@@
While a trained classification model is necessary for using the intended BCI application, the calibration recording delays a deployment and may be prohibitive specifically for users with a limited attention span. In general, the necessity of calibration may impede the acceptance and widespread adoption of BCIs by patients and healthy users. Encouragingly, recent advancements in BCI technology have surfaced which offer potential solutions to alleviate this challenge.
@@
The first advancement involves selecting an informative brain signal feature to minimize the duration of the calibration phase. BCIs can be driven by various brain signals, often evoked by advanced stimulus protocols. Many popular stimulus protocols induce one or more event-related potentials (ERPs). Among these, the visual evoked potential (VEP), triggered by a flash, stands out. VEP-based BCIs are widely embraced due to their effectiveness across diverse user populations [volosyak2020]. The VEP can be effectively used in three different ways [bin2009].
@@
Firstly, in a BCI based on time-modulated VEP (t-VEP), stimuli are sequentially presented to reduce temporal overlap, resulting in a relatively slow paradigm. Secondly, in a frequency-modulated (f-VEP)-based BCI, each stimulus simultaneously and rapidly flashes at a unique frequency and phase [nakanishi2017]. Despite its speed, f-VEP faces limitations due to the restricted range of narrow-band options and potential artefacts that may obscure signals. Thirdly, in a code-modulated (c-VEP)-based BCI, each stimulus rapidly flashes with a pseudo-random noise-code [martinez2021]. In this protocol, stimulus sequences are optimized to be dissimilar, ensuring that their corresponding brain activity is dissimilar as well. Remarkably, c-VEP BCI has recently demonstrated unprecedented performance [shi2024].
@@
The second advancement involves the choice of the decoding approach. In the case of c-VEP, a method was developed, termed ‘reconvolution’, which relies on a forward model embedded in a canonical correlation analysis (CCA) [thielen2015]. This model characterizes the response to a sequence of flashes as the linear summation of responses to individual flashes. This reconvolution approach substantially reduced the number of trainable parameters while it simultaneously increased the number of samples that were available as training data points. This not only decreased the required training data but also empowered the model to predict responses to unseen stimulus sequences. Recently, this reconvolution CCA method was shown to achieve remarkable performances on c-VEP data event without the need for a calibration session, by finding the stimulus sequence that best fits the data in a trial [thielen2021].
@@
In a recent development, a novel classification approach was introduced for ERP-based BCI. The approach, termed ‘unsupervised mean-difference maximization’ (UMM) does not require labeled calibration data [sosulski2023]. UMM does not act on every single epoch, i.e., the evoked response of a single stimulus, but on a set of epochs belonging to one control command, e.g., one trial that leads to the selection of a symbol in a spelling application. Therefore, UMM has similarities to the aforementioned CCA method. While the objective of CCA is to maximize the explained variance, UMM identifies an attended target symbol by maximizing the distance between target and non-target ERPs belonging to one trial. Diverging from the CCA method, UMM incorporates several regularization approaches, including the use of block-Toeplitz covariance matrices [sosulski2022] for determining domain-specific distances, and it can take advantage of a built-in confidence metric. Collectively, UMM has demonstrated impressive performance across various ERP datasets, without the need for a training session [sosulski2023].
@@
Dataset: We assessed the efficacy of the CCA and UMM zero-training approaches using an open-access c-VEP dataset [thielen2023dataset]. Comprehensive details about this dataset can be found in the original study [thielen2021]. For the current study, we only used the part of this dataset labeled as ‘offline experiment’, in which 30 participants engaged in a copy-spelling task. EEG data were recorded from 8 electrodes placed following the 10-10 system (Fz, T7, O1, POz, Oz, Iz, O2, T8) amplified using a Biosemi Active2 amplifier and sampled at a frequency of 512 Hz.
Throughout the experiment, participants interacted with a $4\times 5$ matrix speller displayed on a 12.9 in iPad Pro with a 60 Hz refresh rate and a $1920\times 1080$ px resolution. The $N=20$ cells within this matrix each measured 3.1 cm $\times$ 2.8 cm, with a 0.4 cm separation both horizontally and vertically between cells. The cells were presented against a mean-luminance gray background. Each cell $i\in\{1,\dots,N\}$ was luminance modulated using a unique binary stimulus sequence at full contrast, with a 1 encoding a white cell and a 0 a black cell. The stimulus sequences were carefully chosen from an optimized subset of Gold codes [gold1967]. These sequences were modulated such that they contained flashes of only two durations: a short flash of 16.67 ms and a long flash of 33.33 ms. The sequences had a length of $126=2*(2^{6}-1)$ bits and at 60 Hz cycling through a stimulation code once took 2.1 s.
@@
Alternatively to this instantaneous CCA, CCA can learn across trials. Specifically, previous trials can be included to improve the estimates for the current trial, as described in [thielen2021]. Equation 1 can be formulated using the spatio-temporal cross-covariance ${\bm{\Sigma}}_{\mathbf{X}\mathbf{M}_{i}}\in\mathbb{R}^{C\times M}$, the spatial covariance ${\bm{\Sigma}}_{\mathbf{X}}\in\mathbb{R}^{C\times C}$ and the temporal covariance ${\bm{\Sigma}}_{\mathbf{M}_{i}}$:
@@
In summary, in this work, we applied two versions of CCA. For both we used the empirical covariance matrix identical to the original work [thielen2021]. The first version, denoted CCA_e1, was instantaneous and estimated the covariance from the current trial only. The second version, denoted CCA_ec, was cumulative and used previous trials for covariance estimation to facilitate decoding of the current trial. Code for the CCA approach is available at https://github.com/thijor/pyntbci.
@@
Instead of determining the attended symbol of the current trial by maximizing the Euclidean distance between flash and non-flash ERPs across all $N$-many hypotheses, the metric is first normalized using the inverse of the global feature covariance matrix ${\bm{\Sigma}}\in\mathbb{R}^{D\times D}$ to better cope with non-spherical feature distributions in the feature space $\mathbb{R}^{D}$, known as the Mahalanobis distance:
@@
Please note that the covariance matrix can be estimated based on the epochs of the current trial only, which makes UMM an instantaneous decoding approach that does not require calibration data. Due to the challenging ratio of the feature dimensionality and the number of epochs contained in a single trial, we used a block-Toeplitz regularization with tapering to obtain a more robust estimate of the covariance matrix [sosulski2022].
Alternatively to this instantaneous use of UMM, knowledge from previous trials about both, the estimated class means and the covariance matrix can be included for obtaining improved estimates for the current trial, as described in [sosulski2023]. Specifically, as the covariance matrix can be calculated without label information the covariance matrix can more robustly be estimated by also using epochs from previous trials [sosulski2023]. A similar approach can be used to more robustly estimate the flash and non-flash ERPs, by using information from previous trials. However, for this, label information is required. UMM simply uses its own predictions from previous trials as pseudo labels (sometimes also referred to as naive labeling). This approach is made more robust, by weighting the mean estimates from previous trials by UMM’s confidence in these previous trials [sosulski2023]. Specifically, if UMM is very certain of its own prediction, these ERP means will have more weight in later trials and vice versa.
Figure 1: Bandpass hyper-parameters for CCA and UMM. Depicted are the grand average classification accuracy for CCA and UMM across varying highpass (left) and lowpass (right) cutoff values. Here, a single-trial duration of 31.5 s is used. When varying the highpass, the lowpass remained at 40 Hz, and when varying the lowpass, the highpass remained at 6 Hz. In order, the symbols behind a method refer to: the type of covariance matrix being empirical (e) or block-Toeplitz (t); covariance matrices computed instantaneously (1) or cumulative (c); and the mean vectors (of UMM) computed either instantaneously (1) or using a weighted cumulative average (w). The dashed gray line denotes the theoretical chance level (5%).
In summary, for this work, we used two versions of UMM. For both versions we used the block-Toeplitz regularized covariance matrix. The first version, denoted UMM_t11, is instantaneous by estimating the covariance and means only from the current trial. The second version, denoted UMM_tcw, is cumulative by using previous trials to facilitate decoding of the current trial. For an overview of all method’s abbreviations used, please refer to the legend of Fig. 1. Code for the UMM approach is available at: https://github.com/jsosulski/umm_demo.
@@ RESULTS
Figure 2: Decoding curve for CCA and UMM. Depicted are the grand average classification accuracy for CCA and UMM across varying single-trial durations. Here, a bandpass of 6-50 Hz is used. For a definition of method names, see 2. The dashed gray line denotes the theoretical chance level (5%).
@@
The instantaneous CCA method (CCA_e1) achieved a performance of 0.06 at 1.05 s, while the instantaneous UMM (UMM_t11) reached 0.09. At this time point, UMM_t11 significantly outperformed CCA_e1. Instead, at any further time point CCA_e1 significantly surpassed the accuracy of UMM_t11. At the maximum trial length of 31.5 s, CCA_e1 reached a classification of 0.96 which was significantly higher than UMM_t11 with 0.89.
The cumulative versions always outperformed the instantaneous version for both CCA and UMM. Notably, the instantaneous CCA (CCA_e1) performed almost on par with its cumulative version (CCA_ec) for trials longer than 25.2 s. Specifically, at 27.3 s, CCA_e1 reached an accuracy of 0.95 while CCA_ec reached 0.97, which was not significantly higher. Also at the 29.4 and 31.5 s trials, CCA_ec did not significantly surpass CCA_e1.
@@ DISCUSSION
We introduced UMM to c-VEP BCI, a calibration-free method originating from the ERP domain. We conducted a systematic offline comparison with the conventional c-VEP method employing CCA. Both methods underwent evaluation in an instantaneous manner, classifying each single trial without prior knowledge, as well as in a cumulative way, utilizing previously classified trials as training data. The ultimate goal was to establish a c-VEP BCI with enhanced usability and broader adoption potential by eliminating the initial calibration session. We gave both methods access to EEG data only that would be available in an online experiment such that we don’t expect a major problem for online use. For CCA approaches, online performance was reported in [thielen2021] where the non-instantaneous version matched a supervised model after a warm-up period.
Both the conventional method, CCA, and the novel approach, UMM, hinge on the principle that selecting a single symbol from a set of candidate symbols is considerably simpler than the reconstructing the entire stimulus sequence. Using 20 symbols in this study, one has to evaluate 20 candidate stimulus sequences only for a symbol selection. In contrast, the exhaustive reconstruction of the stimulus would invole an exponential growth, reaching $2^{60}$ potential sequences for a 1-second stimulus at a presentation rate of 60 Hz.
@@
Furthermore, we demonstrated the capability of instantaneous classification, where CCA reached higher than 90 % accuracy at 14.70-second trials and UMM 89 % at 29.40 s. Instantaneous decoding does not learn from, or consider in any way, previous trials. Under stationary conditions, this scarcity of data may not allow the model to reach peak performance. On the other hand, non-stationary feature distributions like latency- or amplitude changes of ERP components over time, will likely not affect instantaneous decoding, while models mainly trained on data collected before the appearance of such feature drifts may suffer from performance degradation.
Both CCA and UMM make different assumptions and exploit them. These assumptions may be met by different datasets to various degrees. For instance, CCA strongly leverages the sequential structure and large overlap between responses to adjacent stimuli in c-VEP datasets. Additionally, the event definition used focusses on the target ERP while discerning between ERPs associated with short and long flashes, but more and different event types can easily be implemented. Furthermore, UMM maintains its operations within the original EEG feature space, whereas CCA operates in the component space.
In contrast, UMM searches for the stimulus sequence with the largest target to non-target ERP distance, potentially rendering it less susceptible to slow drifts in the data, as evidenced by the highpass analysis. Besides, CCA uses the empirical covariance matrix, which can be challenging to estimate with limited data, while UMM employs domain-specific regularization techniques such as shrinkage and a block-Toeplitz covariance matrix [sosulski2022]. Lastly, in the non-instantaneous UMM formulation, ERP mean estimates are improved using previously classified trials, by carefully weighting mean updates based on the confidence of each previous trial. Gaining a comprehensive understanding of the strengths and limitations of both methods bears the potential to develop refined versions tailored to specific characteristics of novel datasets through thoughtful hyper-parameterization.
Essentially, CCA and UMM necessitate a specific stimulus protocol involving repetitions, and both require knowledge about the precise timing and sequence of stimuli within a single trial (i.e., the selection of one symbol). While such information is typically available in BCI protocols using evoked responses, it may not seamlessly extend to other protocols like those based on sensorimotor rhythms. Moreover, these decoding methods are only applicable for benchmarks if the sequence information is provided, as demonstrated in MOABB [aristimunha2023]. This characteristic classifies the studied CCA and UMM methods as semi-supervised, given that stimulus information is requisite, while label information is not required.